# Optimizing Multi-modality Trackers via Sensitivity-regularized Tuning

### v2

@@ IV-A4 Training Details
DiMP [3] ICCV’19 ResNet 48.5 71.7 45.2 66.2 56.7 69.1
Stark-S [55] ICCV’21 ResNet 51.4 75.1 44.6 61.3 55.7 66.7
OSTrack [58] ECCV’22 OSTrack-256 50.7 73.7 53.4 69.6 64.5 77.0
DropTrack [49] CVPR’23 DropTrack-384 52.0 75.9 54.8 72.4 66.1 79.1
@@
FENet [61] ICCV’21 DiMP 63.3 92.4 50.8 70.4 53.8 66.1
AFNet [60] CVPR’23 DiMP 63.8 92.6 51.8 70.8 52.5 63.8
CEUTrack [42] Arxiv’22 OSTrack-B256 56.7 84.2 55.0 71.6 62.0 76.4
ProTrack [57] ACM MM’22 OSTrack-B256 58.0 85.5 47.1 63.2 66.1 78.4
HRTrack [73] ICCV’23 OSTrack-B256 - - - - 63.2 71.9
ViPT [70] CVPR’23 OSTrack-B256 65.1 93.5 59.2 75.8 67.7 80.6
SDSTrack [17] CVPR’24 OSTrack-B256 60.0 87.2 59.7 76.7 67.9 80.9
UnTrack [51] CVPR’24 OSTrack-B256 - - 58.9 75.6 - -
@@
ViPT [70] CVPR’23 DropTrack-B384 65.6 93.9 60.0 77.4 68.6 81.4
SDSTrack [17] CVPR’24 DropTrack-B384 65.6 94.4 60.6 78.0 69.0 82.2
UnTrack [51] CVPR’24 DropTrack-B384 - - 60.0 77.2 - -
@@
MamTrack [41] CVPR’25 HiViT-B384 66.4 94.2 61.6 79.2 - -
SUTrack [9] AAAI’25 SUTrack-B384 - - 63.1 79.6 - -
@@
TABLE II: Quantitative comparison on the RGB-Depth and RGB-Thermal datasets. The best are marked with “bold”, and the second best are marked with “underline”.
@@
DiMP [3] ICCV’19 ResNet 46.3 42.8 44.5 70.3 73.1 54.3 42.8 54.4 42.1 62.5
Stark-S [55] ICCV’21 ResNet 39.3 37.6 38.4 65.4 62.8 48.2 37.7 46.3 48.9 66.5
OSTrack [58] ECCV’22 OSTrack-B256 53.6 52.2 52.9 80.3 83.3 67.6 47.2 58.5 54.9 72.9
DropTrack [49] CVPR’23 DropTrack-B384 56.4 55.8 56.1 81.5 85.1 69.2 47.7 59.4 57.2 75.8
@@
SPT [71] AAAI’23 Stark-S 52.7 54.9 53.8 79.8 85.1 65.1 33.4 46.7 55.5 78.6
APFNet [53] AAAI’22 Stark-S 51.6 51.4 51.5 79.0 83.7 64.2 36.2 50.0 57.9 82.7
ProTrack [57] ACM MM’22 OSTrack-B256 58.3 57.3 57.8 80.1 80.2 65.1 42.0 53.8 59.9 79.5
ViPT [70] CVPR’23 OSTrack-B256 59.2 59.6 59.4 81.5 87.1 72.1 52.5 65.1 61.7 83.5
SDSTrack [17] CVPR’24 OSTrack-B256 61.9 60.9 61.4 81.2 88.3 72.8 53.1 66.5 62.5 84.8
UnTrack [51] CVPR’24 OSTrack-B256 61.1 60.8 61.0 82.0 86.9 72.1 51.3 63.7 62.5 84.2
@@
ViPT [70] CVPR’23 DropTrack-B384 62.6 61.9 62.2 81.7 87.3 72.3 52.5 65.6 63.4 84.7
SDSTrack [17] CVPR’24 DropTrack-B384 63.3 62.2 62.7 81.4 88.6 73.0 55.1 69.0 65.0 87.1
UnTrack [51] CVPR’24 DropTrack-B384 62.4 61.7 62.0 82.1 87.1 72.2 52.4 65.9 64.1 85.2
@@
MamTrack [41] CVPR’25 HiViT-B384 - - - - - - 54.2 67.4 62.4 84.4
SUTrack [9] AAAI’25 SUTrack-B384 58.8 58.6 58.7 83.0 90.6 75.4 60.8 75.6 69.2 92.1
@@
TABLE III: Quantitative comparison with OneTracker.
@@
OneTracker [16] CVPR’24 OSTrack-B384 60.8 76.7 60.7 60.4 60.9 81.9 87.2 72.7 53.8 67.2 64.2 85.7
@@
IV-B Comparison with State-of-the-Art Methods
Extensive comparative analyses are presented in Table I, Table II and Table III, where our method demonstrates excellent performance on all multi-modal benchmarks after incorporating the proposed regularized tuning strategies. The corresponding precision and success plots are illustrated in Fig. 6, further substantiating the quantitative results. Evidently, we can observe that both the RGB-only and the cross-modal trackers are becoming increasingly profitable with pre-trained models. In particular, cross-modal approaches exhibit substantial performance gains, highlighting the complementarity between RGB and auxiliary data under complex conditions. Crucially, the remarkable improvements achieved by our method suggest the significance and necessity of developing tailored cross-domain transfer techniques for multi-modal object tracking.
@@ IV-B Comparison with State-of-the-Art Methods
Results on RGB-Event Tracking. As shown in Table I and Table III, our method surpasses all state-of-the-art trackers across all RGB-Event datasets, achieving the highest precision scores of $98.2\%$, $81.5\%$ and $85.6\%$ on the FE108, VisEvent, and CoeSot datasets, respectively. In particular, on FE108, a dataset characterized by low-light conditions and heavy reliance on event information, our approach surpasses the previous best by a notable margin: $3.0\%$ in PR and $2.4\%$ in SR over OSTrack-B256. The full fine-tuning approaches (e.g., CEUTrack, MamTrack) yield limited improvements, while parameter-efficient fine-tuning paradigms (e.g., ViPT, SDSTrack) face performance bottlenecks. This stems from a misfit that impedes cross-modal transfer, emphasizing the effectiveness of our regularized tuning.
@@
Results on RGB-Depth Tracking. As depicted in Table II and Table III, our method outperforms all previous state-of-the-art trackers on DepthTrack, obtaining the top performance $67.1\%$ in F-score. Using the pre-trained OSTrack-B256, our method yields substantial improvements: $2.8\%$ in Pr, $4.5\%$ in Re, and $3.7\%$ in F-score. Similarly, when built on the DropTrack-B384 with richer pre-trained knowledge, our method demonstrates superior performance gains, $3.9\%$ in Pr, $4.9\%$ in Re, and $4.4\%$ in F-score. Furthermore, despite no training on the VOT-RGBD2022 dataset, our method still delivers enhanced performance with a $2.3\%$ gain in EAO over existing baselines. This clearly demonstrates the strong generalization and transferability of our approach in cross-modal tracking scenarios.
@@
Results on RGB-Thermal Tracking. As listed in Table II and Table III, our method achieves $70.1\%$ precision and $56.3\%$ success when evaluated on the pre-trained OSTrack-B256. Furthermore, our method effectively unleashes the potential of the DropTrack, yielding substantial improvements of $4.6\%$ in PR, $3.7\%$ in SR. Remarkably, with SUTrack as our pre-trained baseline, our method sets a new state-of-the-art, reaching $77.8\%$ in PR and $62.9\%$ in SR. Beyond the established benchmarks, our method also excels on the unseen RGBT234 dataset, attaining $93.3\%$ MPR and $70.3\%$ MSR. These results not only underscore the superior performance of our method but also validate its exceptional cross-dataset generalization capabilities.
@@
Attribute Analysis. To comprehensively analyze our method, we present a detailed per-attribute comparison in Fig. 7. Our approach consistently achieves the best performance across nearly all challenging attributes while significantly leading. Specifically, for motion-related sequences from the VisEvent and CoeSot datasets, our method delivers the best results, highlighting strong robustness against motion-induced degradation. Particularly, it yields precision gains of $5.6\%$ under Motion Blur, $3.0\%$ under Camera Motion. On the LasHeR, our regularization strategy yields notable improvements under extreme lighting conditions: $13.4\%$ under Illumination Variation, $6.2\%$ under Low Illumination, and $5.3\%$ under Over Exposure. Moreover, our method also maintains superior performance across other challenging attributes such as Out-of-View and Frame Lost.
IV-C Ablation Study and Analysis
@@
We conducted ablative analyses to verify the effectiveness and characteristics of sensitivity-aware regularization tuning. In the ablation section, “full fine-tuning” refers to training the entire backbone, excluding the box head (as Section III-A). All methods use the pre-trained OSTrack-B256 weights unless otherwise specified.
@@
Effectiveness of Proposed Sensitivities. We conduct a series of comprehensive experiments to uncover the interplay and effectiveness of the two proposed sensitivity items, as summarized in Table IV. A comparison of (a) and (b) demonstrates that fine-tuning significantly enhances the domain adaptation ability. Further, the contrast between (b) and (c) (or (d) and (e)) figures out that the prior sensitivity regularization scheme significantly improves SR by $2.4\%$ and PR by $3.2\%$ on LasHeR, highlighting its effectiveness. Similarly, the comparison between (b) and (d) (or (c) and (e)) reveals that rebalancing gradients optimized with diverse transfer data leads to substantial improvements. Notably, the simultaneous application of both prior and transfer sensitivity yields significantly greater improvements than either method alone, proving their complementary nature. While the transfer sensitivity penalty alone yields only modest improvement on FE108, it still contributes to the overall leading performance of our method.
@@
TABLE IV: Ablative study results of the proposed key components. “FFT” refers to full fine-tuning of the backbone; “PS” represents prior sensitivity; “TS” is transfer sensitivity. (a) is the zero-shot performance; (b) serves as our baselines; (e) denotes the complete regularized tuning framework.
@@
Figure 8: Operation-wise eigenvalue spectra for measuring prior sensitivity. Eigenvalues of each operation in OSTrack-B256, obtained during noise probing, are sorted in ascending order and plotted on a logarithmic scale. For each operation, the estimated large eigenvalues show a clear tendency to stabilize, indicating convergent behavior toward the true maximum eigenvalues.
Eigenvalue Budget (Top-$K$) for Prior Sensitivity. In this section, we examine the effect of the eigenvalue count $K$ in Eq. (8). For each operation, we record up to the 100 largest eigenvalues during Rayleigh-quotient probing and compute the prior sensitivity using the top-$K$ of them. As illustrated in Fig. 8, the eigenvalue spectra exhibit a clear saturation behavior in the high-rank regime, indicating convergent behavior toward the true maximum eigenvalues. Consistent with this observation, Table V shows that decreasing $K$ from 20 to 1 has little impact: $K=5$ performs on par with $K=10$ (the best), and even $K=1$ still yields a clear improvement over the baseline. This indicates that the effectiveness of our regularizer depends primarily on the relative ranking of sensitivities rather than their absolute magnitudes. Under this approximation, the key challenge lies in accurately identifying the operations most sensitive to the generalization gap.
TABLE V: Ablation analysis of low-dimension eigenvalue count $K$ on prior sensitivity estimation.
@@
Sensitivity Harmony Coefficient $\kappa$. In this section, we evaluate the impact of different sensitivity harmony coefficients, $\kappa$ as defined in Eq. (14), with the detailed results presented in Table VI. By systematically increasing $\kappa$ from 0.1 to 0.9, we observe that a moderately large coefficient (e.g., 0.6) yields the best performance, outperforming both smaller (e.g., 0.3) and larger values (e.g., 0.9). This finding suggests that while both prior sensitivity and transfer sensitivity contribute to performance improvements from different perspectives, a delicate balance between them further optimizes the fine-tuning process, promoting better generalization and adaptability.
TABLE VI: Ablation analysis of sensitivity harmony coefficient on the FE108 dataset.
@@
Impact of Box Head Tuning. In our default setting, “full fine-tuning” updates the backbone while keeping the box head frozen. To assess the impact of tuning the box head, we also examine a more aggressive variant, complete fine-tuning, where the box head is updated as well. As shown in Table VII, unfreezing the box head further exacerbates overfitting. While our regularization method remains effective, the results reveal an important caveat: unfreezing the box head during cross-modal adaptation can be risky, as it tends to disrupt the modal-agnostic object association knowledge.
TABLE VII: Ablation results of box head tuning. “Box” denotes tuning the box head, “S-Reg” represents the sensitivity-aware regularization tuning method.
@@
Compatibility with PEFT Methods. Existing PEFT methods [70, 17, 51] typically freeze pre-trained parameters and update only a minimal number of additional parameters, which can hinder sufficient optimization. To evaluate the compatibility of our proposed regularization techniques with these methods, we unfreeze their backbones and retrain them with our regularization applied. As shown in Table VIII, our approach significantly enhances ViPT’s performance on VisEvent, yielding gains of $2.1\%$ in SR and $2.0\%$ in PR. Notably, even for the unified tracker UnTrack, our approach achieves superior results on LasHeR: $2.7\%$ in SR and $3.7\%$ in PR. This demonstrates that overly rigid constraints on pre-trained models limit their transfer potential. However, our method negatively impacts SDSTrack, likely because it optimizes the pre-trained parameters, whereas SDSTrack relies on modal-specific adapters trained from scratch.
TABLE VIII: Compatibility study results of the proposed regularization tuning method on existing PEFT methods. “FFT” denotes full fine-tuning of the backbone, while “S-Reg” represents the sensitivity-aware regularization tuning scheme.
@@
Compatibility with Single-modality Methods. This work aims to mitigate the misfitting issue when adapting the foundation trackers to downstream tasks. A central question is how the proposed regularization methods perform on single-modality data. To investigate this, we conduct ablation studies highlighting their impact on different modalities. As shown in Table IX, both RGB and auxiliary modalities benefit significantly from our regularization techniques. For example, the RGB and depth modalities achieve F-score gains of $5.3\%$ and $4.0\%$, respectively, on the DepthTrack dataset. Despite substantial distribution differences, our method significantly and consistently enhances the adaptability of auxiliary modalities across various tasks. These findings underscore the importance and necessity of applying constraints when transferring the pre-trained trackers to downstream domains.
TABLE IX: Compatibility study results of the proposed regularization tuning method on single-modal data. “S-Reg” represents the regularization tuning method.
@@
Impact of Event Representations. In this work, we focus on constructing suitable frame-like event representations tailored to cross-modal transfer. In this experiment, we keep the same training setup, varying only the input event representations. As shown in Table X, surface-based representations (TSLTD, Time Surface) perform poorly, likely due to the randomness of event timestamps. In contrast, count-based representations (Event Count, Event Frame) offer more robust performance. The time-interpolation-based Event Volume further improves results by utilizing time-weighted and multi-channel methods to preserve spatio-temporal information. Notably, we employ simple color-polarity event images (with ViPT) to align with the RGB pre-trained model, leading to superior performance.
TABLE X: Ablation results of different event representations.

### v3

@@ 4.1.1 Benchmark Datasets
DiMP (Bhat et al., 2019) ICCV’19 ResNet 48.5 71.7 45.2 66.2 56.7 69.1
Stark-S (Yan et al., 2021a) ICCV’21 ResNet 51.4 75.1 44.6 61.3 55.7 66.7
OSTrack (Ye et al., 2022) ECCV’22 OSTrack-256 50.7 73.7 53.4 69.6 64.5 77.0
DropTrack (Wu et al., 2023) CVPR’23 DropTrack-384 52.0 75.9 54.8 72.4 66.1 79.1
@@
FENet (Zhang et al., 2021a) ICCV’21 DiMP 63.3 92.4 50.8 70.4 53.8 66.1
AFNet (Zhang et al., 2023a) CVPR’23 DiMP 63.8 92.6 51.8 70.8 52.5 63.8
CEUTrack (Tang et al., 2022) Arxiv’22 OSTrack-B256 56.7 84.2 55.0 71.6 62.0 76.4
ProTrack (Yang et al., 2022) ACM MM’22 OSTrack-B256 58.0 85.5 47.1 63.2 66.1 78.4
HRTrack (Zhu et al., 2023c) ICCV’23 OSTrack-B256 - - - - 63.2 71.9
ViPT (Zhu et al., 2023a) CVPR’23 OSTrack-B256 65.1 93.5 59.2 75.8 67.7 80.6
SDSTrack (Hou et al., 2024) CVPR’24 OSTrack-B256 60.0 87.2 59.7 76.7 67.9 80.9
UnTrack (Wu et al., 2024b) CVPR’24 OSTrack-B256 - - 58.9 75.6 - -
@@
ViPT (Zhu et al., 2023a) CVPR’23 DropTrack-B384 65.6 93.9 60.0 77.4 68.6 81.4
SDSTrack (Hou et al., 2024) CVPR’24 DropTrack-B384 65.6 94.4 60.6 78.0 69.0 82.2
UnTrack (Wu et al., 2024b) CVPR’24 DropTrack-B384 - - 60.0 77.2 - -
@@
MamTrack (Sun et al., 2025) CVPR’25 HiViT-B384 66.4 94.2 61.6 79.2 - -
SUTrack (Chen et al., 2025b) AAAI’25 SUTrack-B384 - - 63.1 79.6 - -
@@
4.1.2 Evaluation Metrics
Adhering to recognized standards (Hou et al., 2024; Wu et al., 2024b; Hong et al., 2024; Sun et al., 2025), we evaluated tracking performance with the following metrics. For all RGB-Event benchmarks (FE108, VisEvent and CoeSot), We utilize two widely used metrics, i.e., success rate (SR), precision rate (PR). For DepthTrack, precision (Pr) and recall (Re) are used, with F-score ($F=\frac{2Re\cdot Pr}{Re+Pr}$) as the primary measure. For VOT-RGBD2022, we assess accuracy (Acc), robustness (Rob), and expected average overlap (EAO). For LasHeR, success rate (SR) and precision rate (PR) are employed. For RGBT234, maximum success rate (MSR) and maximum precision rate (MPR) served as evaluation metrics. Notably, our SR/PR precisely aligns with SR/PR in ViPT and OneTrack, Suc/Pre in SDSTrack, Success/Precision in UnTrack, or AUC/P in SUTrack. For all tracking metrics, the larger, the better.
4.1.3 Pre-trained Models and Compared Methods
In this work, we choose three prototypical one-stream RGB-based trackers as pre-trained baselines: OSTrack (Ye et al., 2022), the most widely adopted model; DropTrack (Wu et al., 2023) and SUTrack (Chen et al., 2025b), which offer improved generalization and performance. Notably, OSTrack and DropTrack are built upon ViT-B/16 (Dosovitskiy et al., 2020), while SUTrack adopts the HiViT-B/16 (Zhang et al., 2023b) as the backbone. Following their pre-trained configurations, these models differ in input resolution: OSTrack ($\times128$$\times256$), DropTrack ($\times192$$\times384$), and SUTrack ($\times192$$\times384$) for template and search regions, respectively. To comprehensively validate the effectiveness of our method, we conduct the following experiments in Table 1 and Table 2. First, we construct strong RGB-based single-modal baselines under full fine-tuning. Then, we evaluate our method under both full and parameter-efficient fine-tuning paradigms. For fair comparison, all methods are grouped according to the pre-trained trackers. Notably, XTrack (Tan et al., 2025b), OneTracker (Hong et al., 2024) and MamTrack (Sun et al., 2025) adopt non-standard experimental settings and lack open-source implementations. XTrack (Tan et al., 2025b) and OneTracker use a $\times384$ pre-trained OSTrack, incompatible with standard benchmarks, so we compare it separately in Table 3. MamTrack, with unclear pretraining details but sharing the HiViT-B/16 backbone with SUTrack, is included in SUTrack-based comparisons.
4.1.4 Training Details
We follow the data processing pipeline of SDSTrack (Hou et al., 2024) across all datasets, converting event data into color-polar event images, with no preprocessing for other auxiliary modalities. The models are trained on 8 NVIDIA 3090Ti GPUs with a batch size of 192 and 30 epochs. Each epoch involves sampling 80k samples. We utilize the AdamW optimizer with a learning rate set to $1\times 10^{-4}$ and a weight decay set to $10^{-4}$.
4.2 Comparison with State-of-the-Art Methods
Extensive comparative analyses are presented in Table 1, Table 2 and Table 3, where our method demonstrates excellent performance on all multi-modal tracking benchmarks after incorporating the proposed regularized tuning strategies. The corresponding precision and success plots are illustrated in Fig. 6, further substantiating the quantitative results. Evidently, we can observe that both the RGB-only and the cross-modal trackers are becoming increasingly profitable with pre-trained models. In particular, cross-modal approaches exhibit substantial performance gains, highlighting the complementarity between RGB and auxiliary data under complex conditions. Crucially, the remarkable improvements achieved by our method suggest the significance and necessity of developing tailored cross-domain transfer techniques for multi-modal object tracking.
Table 2: Quantitative comparison on the RGB-Depth and RGB-Thermal datasets. The best are marked with “bold”, and the second best are marked with “underline”.
@@ 4.2 Comparison with State-of-the-Art Methods
DiMP (Bhat et al., 2019) ICCV’19 ResNet 46.3 42.8 44.5 70.3 73.1 54.3 42.8 54.4 42.1 62.5
Stark-S (Yan et al., 2021a) ICCV’21 ResNet 39.3 37.6 38.4 65.4 62.8 48.2 37.7 46.3 48.9 66.5
OSTrack (Ye et al., 2022) ECCV’22 OSTrack-B256 53.6 52.2 52.9 80.3 83.3 67.6 47.2 58.5 54.9 72.9
DropTrack (Wu et al., 2023) CVPR’23 DropTrack-B384 56.4 55.8 56.1 81.5 85.1 69.2 47.7 59.4 57.2 75.8
@@
SPT (Zhu et al., 2023b) AAAI’23 Stark-S 52.7 54.9 53.8 79.8 85.1 65.1 33.4 46.7 55.5 78.6
APFNet (Xiao et al., 2022) AAAI’22 Stark-S 51.6 51.4 51.5 79.0 83.7 64.2 36.2 50.0 57.9 82.7
ProTrack (Yang et al., 2022) ACM MM’22 OSTrack-B256 58.3 57.3 57.8 80.1 80.2 65.1 42.0 53.8 59.9 79.5
ViPT (Zhu et al., 2023a) CVPR’23 OSTrack-B256 59.2 59.6 59.4 81.5 87.1 72.1 52.5 65.1 61.7 83.5
SDSTrack (Hou et al., 2024) CVPR’24 OSTrack-B256 61.9 60.9 61.4 81.2 88.3 72.8 53.1 66.5 62.5 84.8
UnTrack (Wu et al., 2024b) CVPR’24 OSTrack-B256 61.1 60.8 61.0 82.0 86.9 72.1 51.3 63.7 62.5 84.2
@@
ViPT (Zhu et al., 2023a) CVPR’23 DropTrack-B384 62.6 61.9 62.2 81.7 87.3 72.3 52.5 65.6 63.4 84.7
SDSTrack (Hou et al., 2024) CVPR’24 DropTrack-B384 63.3 62.2 62.7 81.4 88.6 73.0 55.1 69.0 65.0 87.1
UnTrack (Wu et al., 2024b) CVPR’24 DropTrack-B384 62.4 61.7 62.0 82.1 87.1 72.2 52.4 65.9 64.1 85.2
@@
MamTrack (Sun et al., 2025) CVPR’25 HiViT-B384 - - - - - - 54.2 67.4 62.4 84.4
SUTrack (Chen et al., 2025b) AAAI’25 SUTrack-B384 58.8 58.6 58.7 83.0 90.6 75.4 60.8 75.6 69.2 92.1
@@
Table 3: Quantitative comparison with OneTracker and XTrack.
@@
OneTracker (Hong et al., 2024) CVPR’24 OSTrack-B384 60.8 76.7 60.7 60.4 60.9 81.9 87.2 72.7 53.8 67.2 64.2 85.7
XTrack (Tan et al., 2025b) ICCV’25 OSTrack-B384 60.9 77.5 61.8 62.0 61.5 82.1 88.8 74.0 55.7 69.1 64.9 87.4
@@
Results on RGB-Event Tracking. As shown in Table 1 and Table 3, our method surpasses all state-of-the-art trackers across all RGB-Event datasets, achieving the highest precision scores of $98.2\%$, $81.5\%$ and $85.6\%$ on the FE108, VisEvent, and CoeSot datasets, respectively. In particular, on FE108, a dataset characterized by low-light conditions and heavy reliance on event information, our approach surpasses the previous best by a notable margin: $3.0\%$ in PR and $2.4\%$ in SR over OSTrack-B256. The full fine-tuning approaches (e.g., CEUTrack, MamTrack) yield limited improvements, while parameter-efficient fine-tuning paradigms (e.g., ViPT, SDSTrack) face performance bottlenecks. This stems from a misfit that impedes cross-modal transfer, emphasizing the effectiveness of our regularized tuning.
@@
Results on RGB-Depth Tracking. As depicted in Table 2 and Table 3, our method outperforms all previous state-of-the-art trackers on DepthTrack, obtaining the top performance $67.1\%$ in F-score. Using the pre-trained OSTrack-B256, our method yields substantial improvements: $2.8\%$ in Pr, $4.5\%$ in Re, and $3.7\%$ in F-score. Similarly, when built on the DropTrack-B384 with richer pre-trained knowledge, our method demonstrates superior performance gains, $3.9\%$ in Pr, $4.9\%$ in Re, and $4.4\%$ in F-score. Furthermore, despite no training on the VOT-RGBD2022 dataset, our method still delivers enhanced performance with a $2.3\%$ gain in EAO over existing baselines. This clearly demonstrates the strong transferability of our approach in cross-modal tracking scenarios.
@@
Results on RGB-Thermal Tracking. As listed in Table 2 and Table 3, our method achieves $70.1\%$ precision and $56.3\%$ success when evaluated on the pre-trained OSTrack-B256. Furthermore, our method effectively unleashes the potential of the DropTrack, yielding substantial improvements of $4.6\%$ in PR, $3.7\%$ in SR. Remarkably, with SUTrack as our pre-trained baseline, our method sets a new state-of-the-art, reaching $77.8\%$ in PR and $62.9\%$ in SR. Beyond the established benchmarks, our method also excels on the unseen RGBT234 dataset, attaining $93.3\%$ MPR and $70.3\%$ MSR. These results not only underscore the superior performance of our method but also validate its exceptional cross-dataset generalization capabilities.
@@
Attribute Analysis. To comprehensively analyze our method, we present a detailed per-attribute comparison in Fig. 7. Our approach consistently achieves the best performance across nearly all challenging attributes while significantly leading. Specifically, for motion-related sequences from the VisEvent and CoeSot datasets, our method delivers the best results, highlighting strong robustness against motion-induced degradation. Particularly, it yields precision gains of $5.6\%$ under Motion Blur, $3.0\%$ under Camera Motion. On the LasHeR, our regularization strategy yields notable improvements under extreme lighting conditions: $13.4\%$ under Illumination Variation, $6.2\%$ under Low Illumination, and $5.3\%$ under Over Exposure. Moreover, our method also maintains superior performance across other challenging attributes such as Out-of-View and Frame Lost. Overall, our models performs substantially better across diverse challenging conditions.
4.3 Ablation Study and Analysis
@@
We conducted ablative analyses to verify the effectiveness and characteristics of significance-aware regularization tuning. In the ablation section, “full fine-tuning” refers to training the entire backbone, excluding the box head (as in Section 3.1). All methods use the pre-trained OSTrack-B256 weights unless specified otherwise.
@@
Effectiveness of Parameter Significance. We conduct a series of comprehensive experiments to uncover the interplay and effectiveness of the two proposed parameter significance items, as summarized in Table 4. Comparing (a) and (b) demonstrates that fine-tuning significantly enhances the domain adaptation ability. Further, the contrast between (b) and (c) (or (d) and (e)) figures out that the prior significance regularization significantly improves SR by $2.4\%$ and PR by $3.2\%$ on LasHeR, highlighting its effectiveness. Similarly, comparing (b) and (d) (or (c) and (e)) reveals that rebalancing gradients optimized with diverse transfer data leads to substantial improvements. Notably, the simultaneous application of both prior and transfer significance yields greater improvements than either method alone, proving their complementary nature. While the transfer significance penalty alone yields only modest improvement on FE108, it still contributes to the overall leading performance of our method.
@@
Table 4: Ablative study results of the proposed key components. “FFT” refers to full fine-tuning of the backbone; “PS” represents prior significance; “TS” is transfer significance. (a) is the zero-shot performance; (b) serves as our baselines; (e) denotes the complete regularized tuning framework.
@@
Table 5: Ablation analysis of low-dimension eigenvalue count $K$ on prior significance estimation.
@@
Figure 8: Operation-wise eigenvalue spectra for measuring prior significance. Eigenvalues of each operation in OSTrack-B256, obtained during noise probing, are sorted in ascending order and plotted on a logarithmic scale. For each operation, the estimated large eigenvalues show a clear tendency to stabilize, indicating convergent behavior toward the true maximum eigenvalues.
Eigenvalue Budget (Top-$K$) for Prior Significance. In this section, we examine the effect of the eigenvalue count $K$ in Eq. (8). For each operation, we record up to the 100 largest eigenvalues during Rayleigh-quotient probing and compute the prior significance using the top-$K$ of them. As illustrated in Fig. 8, the eigenvalue spectra exhibit a clear saturation behavior in the high-rank regime, indicating convergent behavior toward the true maximum eigenvalues. Consistent with this observation, Table 5 shows that decreasing $K$ from 20 to 1 has little impact: $K=5$ performs on par with $K=10$ (the best), and even $K=1$ still yields a clear improvement over the baseline. This indicates that the effectiveness of our regularizer depends primarily on the relative ranking of significance rather than their absolute magnitudes. Under this approximation, the key challenge lies in accurately identifying the operations most sensitive to the generalization gap.
Significance Harmony Coefficient $\kappa$. In this section, we evaluate the impact of different significance harmony coefficients, $\kappa$ as defined in Eq. (14), with the detailed results presented in Table 6. By systematically increasing $\kappa$ from 0.1 to 0.9, we observe that a moderately large coefficient (e.g., 0.6) yields the best performance, outperforming both smaller (e.g., 0.3) and larger values (e.g., 0.9). This finding suggests that while both prior significance and transfer significance contribute to performance improvements from different perspectives, a delicate balance between them further optimizes the fine-tuning process, promoting better generalization and adaptability. Additional significance fusion scheduling strategies are provided in Appendix C.
Table 6: Ablation analysis of significance harmony coefficient on the FE108 dataset.
@@
Impact of Box Head Tuning. In our default setting, “full fine-tuning” updates the backbone while keeping the box head frozen. To assess the impact of tuning the box head, we also examine a more aggressive variant, complete fine-tuning, where the box head is updated as well. As shown in Table 7, unfreezing the box head further exacerbates overfitting. While our regularization method remains effective, the results reveal an important caveat: unfreezing the box head during cross-modal adaptation can be risky, as it tends to disrupt the modal-agnostic object association knowledge.
Table 7: Ablation results of box head tuning. “Box” denotes tuning the box head, “S-Reg” represents the significance-aware regularization tuning method.
@@
Compatibility with PEFT Methods. Existing PEFT methods Zhu et al. (2023a); Hou et al. (2024); Wu et al. (2024b) typically freeze pre-trained parameters and update only a minimal number of additional parameters, which can hinder sufficient optimization. To evaluate the compatibility of our proposed regularization techniques with these methods, we unfreeze their backbones and retrain them with our regularization applied. As shown in Table 8, our approach significantly enhances ViPT’s performance on VisEvent, yielding gains of $2.1\%$ in SR and $2.0\%$ in PR. Notably, even for the unified tracker UnTrack, our approach achieves superior results on LasHeR: $2.7\%$ in SR and $3.7\%$ in PR. This demonstrates that overly rigid constraints on pre-trained models limit their transfer potential. However, our method negatively impacts SDSTrack, likely because it optimizes the pre-trained parameters, whereas SDSTrack relies on modal-specific adapters trained from scratch.
Table 8: Compatibility study results of the proposed regularization tuning method on existing PEFT methods. “FFT” denotes full fine-tuning of the backbone, while “S-Reg” represents the significance-aware regularization tuning scheme.
@@
Compatibility with Single-modality Methods. This work aims to mitigate the misfitting issue when adapting the foundation trackers to downstream tasks. A central question is how the proposed regularization methods perform on single-modality data. To investigate this, we conduct ablation studies highlighting their impact on different modalities. As shown in Table 9, both RGB and auxiliary modalities benefit significantly from our regularization techniques. For example, the RGB and depth modalities achieve F-score gains of $5.3\%$ and $4.0\%$, respectively, on the DepthTrack dataset. Despite substantial distribution differences, our method significantly and consistently enhances the adaptability of auxiliary modalities across various tasks. These findings underscore the importance and necessity of applying constraints when transferring the pre-trained trackers to downstream domains.
Table 9: Compatibility study results of the proposed regularization tuning method on single-modal data. “S-Reg” represents the regularization tuning method.
@@
Impact of Input Event Representations. In this work, we focus on constructing suitable frame-like event representations tailored to cross-modal transfer. In this experiment, we keep the same training setup, varying only the input event representations. As shown in Table 10, surface-based representations (TSLTD, Time Surface) perform poorly, likely due to the randomness of event timestamps. In contrast, count-based representations (Event Count, Event Frame) offer more robust performance. The time-interpolation-based Event Volume further improves results by utilizing time-weighted and multi-channel methods to preserve spatio-temporal information. Notably, we employ simple color-polarity event images (with ViPT) to align with the RGB pre-trained model, leading to superior performance.
Table 10: Ablation results of different event representations.